suppose you have $10,000.00 in cash and you decide to borrow another $10000.00 at 6% interest rate to invest in the stock market. you invest the entire $20,000.00 on an exchange-trade fund with a 11% expected return and a 20% volatility. what is your expected return on your investment?
Answer #1

total return = 11% of 20,000 - 6% of 10,0000 = 1600

expected return on investment = 1600/10000 = 16.00%